a sample of caco3 is50% pure.on heating1.12L of co2(at stp) is obtained.Residue left (assuming non votile impurity) is

Grade:

1 Answers

sunitha mandepudi
11 years ago

1mole of CaCO3 ------------------>1mole of CO2

           ?         <--------------------1.12ltsof CO2

           volume of CaCO3 =(1.12)(22.4)/(22.4)=1.12 lts

its only 50% pure so    

volume of left out residue (50/100)(1.12) = 0.56 lts of impurityis present in the given CaCO3 sample